OpenStudy (anonymous):

Please Help WILL GIVE MEDALS!!!! use the quadratic formula to solve the equation -x^2+7x=8

OpenStudy (studygurl14):

\[x = \frac{ -b \pm \sqrt{b^{2} -4ac} }{ 2a }\] \[ax^{2} + bx + c\] Use this

OpenStudy (anonymous):

can you explain more

OpenStudy (anonymous):

-x^2 - a=-1 7x - b=7 -8 - c=-8

OpenStudy (anonymous):

\[x=\frac{ -(7)\pm \sqrt{(7)^2-4(-1)(-8)} }{ 2(-1) }\]

OpenStudy (anonymous):

do you understand how we got those numbers?

OpenStudy (anonymous):

yes you plugged a,b, and c into the formula

OpenStudy (anonymous):

\[x=\frac{ -7\pm \sqrt{49+32} }{ -2 }\]

OpenStudy (anonymous):

so the answer is \[\frac{7\pm \sqrt{17}}{ 2}\]

OpenStudy (anonymous):

no

OpenStudy (anonymous):

i made a mistake above

OpenStudy (anonymous):

but

OpenStudy (anonymous):

how not

OpenStudy (anonymous):

the 7 and the 2 would be negative

OpenStudy (anonymous):

otherwise it's correct

OpenStudy (anonymous):

ok thank you